A student averages 84 on five tests if the student recieved 84,82,81,and 79 on the first four exams what grade did the student recieve on the fifth exam?
Answer by unlockmath(1688)

Hello,
To find the average of numbers, we need to add all the numbers and divide by 5.
Since one of the numbers is unknown we'll represent by x so the equation will be:
(84+82+81+79+x)/5 = 84 Combine like terms gives us:
(326+x)/5 = 84 Multiply both sides by 5
326 + x = 420 Subtract both sides by 326 gives us:
x = 94 There you go, he needs a 94 on the test to have an average of 84.
RJ Toftness
www.math-unlock.com